OpenStudy (theoreo):

Integral of (4-x^2)^(-3/2) I tried a trig sub, but I must have done it wrong, because my answer isn't even close to the book's.

OpenStudy (mathmath333):

which sub u tried

OpenStudy (theoreo):

I tried \[x = 2\sin(\theta)\] which gave me \[4-4\sin^2 = 4(1-\sin^2) =( 4*\cos^2(\theta))^-3/2\]

OpenStudy (mathmath333):

this sub should work perfectly

OpenStudy (theoreo):

Ok. So if \[(\cos^2)^-3/2 = (\sqrt{\cos^2})^-3= \cos^-3\] right? I got stuck trying to integrate \[\sec^3\] and the answer in my book doesn't involve trig.

OpenStudy (mathmath333):

one cos term should be cancelled and u should have \(\dfrac{\sec^2t}{4}\) right?

OpenStudy (mathmath333):

i mean \(\dfrac{\sec^2\theta}{4}\)

OpenStudy (theoreo):

I don't see how a term would cancel. Could you explain?

OpenStudy (mathmath333):

u have to change \(dx\) to \(d\theta\) so \(x = 2\sin(\theta)\\ \dfrac{dx}{d\theta}=2\cos(\theta)\) had u added that while calculation

OpenStudy (theoreo):

Ah I forgot about that part. Thanks a lot.